# Modeling Temporal Dependencies within the Target for Long-Term Time Series Forecasting

### v2

@@ IV-A1 Datasets
Dataset ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity ILI
Variables 7 7 7 7 21 321 7
Timesteps 17420 17420 69680 69680 52696 26304 966
Granularity 1 hour 1 hour 15 minutes 15 minutes 10 minutes 1 hour 1 week
@@
Model ILI dataset Other datasets
@@
[…]
IV-A2 Baselines
@@
Method MICN +TDTAlign TimesNet +TDTAlign DLinear +TDTAlign
@@
Method PatchTST +TDTAlign iTransformer +TDTAlign SegRNN +TDTAlign
@@ IV-A4 Implementation Details
TABLE V: Statistics of improvements achieved by integrating TDAlign into the baseline method, averaged over seven datasets. Values in parentheses indicate the count of cases showing improvement. For each metric, “Avg improvement” represents the average improvement over six baseline methods, and “Count” indicates the total number of cases showing improvement.
Model MSE MAE $\text{MSE}_{D}$ $\text{MAE}_{D}$ $\rho$
MICN 9.19% (28) 9.72% (28) 9.60% (26) 7.70% (28) 9.03% (28)
TimesNet 6.10% (25) 4.38% (26) 15.78% (27) 11.99% (28) 11.32% (28)
DLinear 7.42% (24) 6.13% (28) 4.57% (18) 4.80% (26) 4.56% (26)
PatchTST 2.58% (21) 3.56% (28) 9.05% (24) 8.71% (28) 9.73% (28)
iTransformer 6.49% (25) 4.95% (28) 11.68% (26) 11.45% (28) 12.17% (28)
SegRNN 1.47% (25) 0.20% (22) 5.72% (24) 3.79% (26) 6.11% (28)
Avg improvement (Count) 5.54% (148) 4.82% (160) 9.40% (145) 8.07% (164) 8.82% (166)
IV-B Main Results and Analysis
@@
Table IV presents a detailed comparison of the forecasting errors (MSE and MAE) between baseline methods and TDAlign. As shown in this table, first, we can clearly observe that our proposed TDAlign framework improves baseline methods by a significant margin in most cases, showcasing its effectiveness and generalizability. Notably, TDAlign improves MICN’s performance across all prediction lengths and datasets. Second, TDAlign contributes to more stable performance, where it shows smaller standard deviations compared to baseline methods in most cases. For instance, DLinear shows non-zero standard deviations across six datasets except for the Electricity dataset, whereas TDAlign demonstrates zero standard deviations on six datasets and exhibits non-zero but significantly smaller standard deviations on the ILI dataset compared to DLinear. Third, implicit learning is insufficient to capture TDT, as demonstrated by SegRNN [22], which maintains sequential order through position embeddings yet still benefits from TDAlign.
TABLE VI: Analysis of the impact of $\mathcal{L}_{D}$ and $\rho$ on the ETTh1 dataset, with the iTransformer [29] as the baseline method. The best performances are highlighted in bold.
@@
96 ① baseline only 0.387±0.001 0.405±0.001
② + $\mathcal{L}_{D}$ 0.377±0.002 0.391±0.000
③ + $\rho$ 0.387±0.001 0.399±0.001
④ + $\mathcal{L}_{D}$ + $\alpha$ 0.378±0.001 0.391±0.001
⑤ + $\mathcal{L}_{D}$ + $\rho$ (TDAlign) 0.375±0.001 0.389±0.000
192 ① baseline only 0.441±0.002 0.436±0.001
② + $\mathcal{L}_{D}$ 0.430±0.001 0.422±0.001
③ + $\rho$ 0.438±0.001 0.428±0.001
④ + $\mathcal{L}_{D}$ + $\alpha$ 0.430±0.002 0.423±0.001
⑤ + $\mathcal{L}_{D}$ + $\rho$ (TDAlign) 0.426±0.001 0.420±0.001
336 ① baseline only 0.491±0.004 0.461±0.003
② + $\mathcal{L}_{D}$ 0.473±0.001 0.445±0.001
③ + $\rho$ 0.484±0.002 0.454±0.002
④ + $\mathcal{L}_{D}$ + $\alpha$ 0.473±0.001 0.445±0.001
⑤ + $\mathcal{L}_{D}$ + $\rho$ (TDAlign) 0.469±0.001 0.442±0.001
720 ① baseline only 0.511±0.010 0.494±0.006
② + $\mathcal{L}_{D}$ 0.470±0.009 0.466±0.007
③ + $\rho$ 0.482±0.010 0.476±0.006
④ + $\mathcal{L}_{D}$ + $\alpha$ 0.471±0.009 0.466±0.007
⑤ + $\mathcal{L}_{D}$ + $\rho$ (TDAlign) 0.460±0.001 0.460±0.001
@@
IV-C Ablation Studies
We conduct ablation experiments on the ETTh1 dataset to evaluate the effectiveness of devised components of TDAlign, using the iTransformer [29] as the baseline method.
IV-C1 Effectiveness of $\mathcal{L}_{D}$ and $\rho$
(a) Using $\tau$-order differencing values to formulate TDT and TDP.
(b) Using change values between time steps separated by the interval $k$ to formulate TDT and TDP.
Figure 4: Ablation of the formulation of temporal dependencies on the ETTh1 dataset, with the iTransformer [29] as the baseline method.
@@ IV-C1 Effectiveness of $\mathcal{L}_{D}$ and $\rho$
  • baseline with $\rho$, where the overall loss is $\mathcal{L}=\rho\cdot\mathcal{L}_{Y}$.
@@
  • baseline with $\mathcal{L}_{D}$ and $\alpha$ (i.e., replacing $\rho$ by a learning weight $\alpha$), where the overall loss is $\mathcal{L}=\alpha\cdot\mathcal{L}_{Y}+(1-\alpha)\cdot\mathcal{L}_{D}$.
  • baseline with TDAlign, where the overall loss is $\mathcal{L}=\rho\cdot\mathcal{L}_{Y}+(1-\rho)\cdot\mathcal{L}_{D}$.
The results are presented in Table VI, from which we observe that: First, TDAlign significantly improves the forecasting performance of iTransformer across all four prediction lengths, as shown by the comparison between and . Second, learning TDT is crucial for improving forecasting performance. Specifically, enables the model to account for the change value between adjacent time steps when forecasting, while enables the model to consider the change direction between adjacent time steps when forecasting. Both settings have superior forecasting performance compared to the baseline method, which overlooks internal correlations across multiple steps.
Figure 5: Loss curves on the validation set of the ETTh1 dataset, with iTransformer [29] as the baseline method.
Third, our proposed new loss function has a more significant impact on TDAlign’s performance than our adaptive loss balancing strategy, where outperform by a large margin. Fourth, our proposed adaptive loss balancing strategy effectively controls the relative importance between the conventional optimization objective and the TDT learning objective. Specifically, by incorporating $\rho$, outperforms , which only incorporates $\mathcal{L}_{D}$, and , where $\rho$ is replaced by a learning weight $\alpha$.
IV-C2 Comparison of the formulation of temporal dependencies
As shown in Figure 4, we evaluate our approach of using first-order differencing values (i.e., change values between adjacent time steps) to formulate temporal dependencies by comparing it with two variants. The first variant uses $\tau$-order differencing values to formulate temporal dependencies. Taking the TDT as an example, its formulation is given by $D=\left\{d^{(\tau)}_{t+1},d^{(\tau)}_{t+2},\cdots,d^{(\tau)}_{t+H}\right\}\in\mathbb{R}^{H\times N}$, where $d^{(\tau)}_{t+i}=d^{(\tau-1)}_{t+i}-d^{(\tau-1)}_{t+i-1}$ and $d^{1}_{t+i}$ corresponds to $d_{t+i}$ defined in Equation 2. We experiment with $\tau\in\{1,2,3,4\}$ The second variant uses change values between time steps separated by the interval $k$ to formulate temporal dependencies. Taking the TDT as an example, its formulation is given by $D=\left\{d_{t+1},d_{t+2},\cdots,d_{t+H}\right\}\in\mathbb{R}^{H\times N}$, where $d_{t+i}=d_{t+i+k}-d_{t+i}$. We experiment with $k\in\{1,6,12,24,48\}$.
@@ IV-C2 Comparison of the formulation of temporal dependencies
The results in Figure 4(a) show that using first-order differencing values (i.e., $\tau=1$) achieves the best performance, with performance degrading as $\tau$ increases. Although higher-order differencing values represent more complex temporal dependencies, they may not provide substantial additional information compared to first-order differencing values. Instead, they tend to overshadow lower-order dependency information and over-constrain the model’s learning, thereby compromising its learning and generalization ability. Moreover, since each $d^{(\tau)}_{t+i}$ involves more time steps, higher-order differencing values become more sensitive to noise, leading to degraded model performance. The results in Figure 4(b) indicate that using change values between adjacent time steps (i.e., $k=1$) yields the best performance. This phenomenon can be attributed to the fact that as $k$ increases, the correlation between $x_{t+i}$ and $x_{t+i+k}$ typically weakens. Learning unreliable temporal dependency information adversely affects model performance. These results demonstrate that using first-order differencing values (i.e., change values between adjacent time steps) to formulate temporal dependencies is simple yet effective.
IV-D Visualization
@@
Figure 6: Comparison of long-term time series forecasting results between our TDAlign and six baseline models.
For an intuitive understanding of the training process of TDAlign, Figure 5 displays the loss curves of the $\text{MSE}_{D}$ and $\rho$ metrics on the ETTh1 dataset’s validation set. We employ iTransformer [29] as the baseline method, with both input and prediction lengths set at 96 steps. These validation loss curves are from experiments with three different settings:
@@
  • baseline with TDAlign, where the overall loss is $\mathcal{L}=\rho\cdot\mathcal{L}_{Y}+(1-\rho)\cdot\mathcal{L}_{D}$.
From these curves, we observe the following: First, the training process of TDAlign is stable, where loss curves of decrease smoothly. Second, explicitly modeling TDT significantly improves the prediction accuracy of both change directions and change values between adjacent time steps, as shown by the lower $\text{MSE}_{D}$ and $\rho$ of compared to . Third, the devised adaptive loss balancing strategy further enhances TDAlign’s performance, as indicated by the further reduction in $\text{MSE}_{D}$ and $\rho$ of compared to .
In order to better illustrate the effectiveness of TDAlign, Figure 6 presents a comprehensive comparison of forecasting results between TDAlign and six baseline methods. Overall, baseline methods produce inaccurate predictions with significant deviations from the ground truth. A closer examination of their forecasting results reveals that these methods struggle to accurately capture both the change values and the change directions. In contrast, TDAlign proposes a TDT learning objective for capturing variation patterns of time series. Moreover, it dynamically adjusts the importance of this TDT learning objective relative to the conventional forecasting objective based on the model’s learning state. Consequently, TDAlign demonstrates remarkable improvements over baseline methods, consistently generating more accurate and realistic forecasting results that align with the ground truth.
V Conclusion

### v3

@@ V-A1 Datasets
Dataset Variables Timesteps Granularity
ETTh1 7 17420 1 hour
ETTh2 7 17420 1 hour
ETTm1 7 69680 15 minutes
ETTm2 7 69680 15 minutes
Weather 21 52696 10 minutes
Electricity 321 26304 1 hour
ILI 7 966 1 week
@@
Method ILI dataset Other datasets
@@
SegRNN 60 $\rightarrow$ {24, 36, 48, 60} 720 $\rightarrow$ {96, 192, 336, 720}
DLinear, PatchTST 104 $\rightarrow$ {24, 36, 48, 60} 336 $\rightarrow$ {96, 192, 336, 720}
iTransformer, MICN, TimesNet 36 $\rightarrow$ {24, 36, 48, 60} 96 $\rightarrow$ {96, 192, 336, 720}
V-A2 Baselines
@@
Method MICN +TDAlign TimesNet +TDAlign DLinear +TDAlign
@@
Method PatchTST +TDAlign iTransformer +TDAlign SegRNN +TDAlign
@@ V-A4 Implementation Details
TABLE V: Averaged performance improvement of TDAlign over 7 datasets, derived from Table IV and Appendix Tables. Values in parentheses indicate the count of cases showing improvement. For each metric, “Avg improvement” represents the average improvement over 6 baseline methods, and “Count” indicates the total number of cases showing a strict improvement (excluding ties).
Method $\Delta\text{MSE}$ $\Delta\text{MAE}$ $\Delta\text{MSE}_{D}$ $\Delta\text{MAE}_{D}$ $\Delta\rho$
MICN +9.19% (28) +9.72% (28) +9.60% (26) +7.70% (28) +9.03% (28)
TimesNet +6.10% (25) +4.38% (26) +15.78% (27) +11.99% (28) +11.32% (28)
DLinear +7.42% (24) +6.13% (28) +4.57% (18) +4.80% (26) +4.56% (26)
PatchTST +2.58% (21) +3.56% (28) +9.05% (24) +8.71% (28) +9.73% (28)
iTransformer +6.49% (25) +4.95% (28) +11.68% (26) +11.45% (28) +12.17% (28)
SegRNN +1.47% (25) +0.20% (22) +5.72% (24) +3.79% (26) +6.11% (28)
Avg improvement (Count) +5.54% (148) +4.82% (160) +9.40% (145) +8.07% (164) +8.82% (166)
TABLE VI: Averaged performance improvement of TDAlign over 6 baseline methods, derived from Table IV. A smaller ADF test statistic indicates a more stationary time series dataset.
Dataset ADF test statistic $\DeltaMSE Δ\DeltaMAE$
Weather -26.68 +2.28% +5.58%
ETTm1 -14.98 +2.13% +3.59%
Electricity -8.44 -0.70% +2.11%
ETTh1 -5.91 +6.29% +5.60%
ETTm2 -5.66 +3.65% +3.92%
ILI -5.33 +6.39% +5.26%
ETTh2 -4.13 +6.85% +5.49%
V-B Main Results and Analysis
@@
Table IV presents a detailed comparison of the forecasting errors (MSE and MAE) between baseline methods and TDAlign. As shown in this table, first, we can clearly observe that our proposed TDAlign framework improves baseline methods by a significant margin in most cases, showcasing its effectiveness and generalizability. Notably, TDAlign improves MICN’s performance across all prediction lengths and datasets. Second, TDAlign contributes to more stable performance, where it shows smaller standard deviations compared to baseline methods in most cases. For instance, DLinear shows non-zero standard deviations across six datasets except for the Electricity dataset, whereas TDAlign demonstrates zero standard deviations on six datasets and exhibits non-zero but significantly smaller standard deviations on the ILI dataset compared to DLinear. Third, implicit learning is insufficient to capture TDT, as demonstrated by SegRNN, which maintains sequential order through position embeddings yet still benefits from TDAlign.
TABLE VII: Analysis of the impact of $\mathcal{L}_{D}$ and $\rho$ on the ETTh1 dataset, with the iTransformer as the baseline method. The best performances are highlighted in bold.
@@
[…]
@@
Table VI presents the average performance improvement conferred by TDAlign across six baseline methods for each dataset. As shown in the table, TDAlign demonstrates consistent performance improvements across all datasets, with particularly pronounced gains on non-stationary datasets. These datasets (e.g., ETTh1, ETTh2, ETTm2, and ILI) exhibit non-stationary characteristics, as confirmed by the Augmented Dickey-Fuller (ADF) test [54]. These series, characterized by time-varying statistical properties, pose a significant challenge as their underlying dynamics are constantly evolving. This is precisely where TDAlign offers a crucial advantage. By explicitly constraining the model to capture local step-wise changes, it equips the baseline methods with the capability to better adapt to these non-stationarities, leading to substantial performance improvements. Specifically, averaged over all experimental prediction lengths and baseline methods, TDAlign achieves MSE and MAE reductions of 6.85% and 5.49% on ETTh2, 6.29% and 5.60% on ETTh1, 3.65% and 3.92% on ETTm2, and 6.39% and 5.26% on ILI, respectively.
V-C Ablation Studies
We conduct ablation experiments on the ETTh1 dataset to assess the effectiveness of devised components in TDAlign and its computational efficiency. All experiments use iTransformer as the baseline method and are repeated five times with different random seeds.
V-C1 Effectiveness of $\mathcal{L}_{D}$ and $\rho$
(a) Using $\tau$-order differencing values.
(b) Using change values between time steps separated by the interval $k$.
(c) Using frequency domain features, i.e. FreDF [51].
Figure 4: Ablation of temporal dependency formulations on the ETTh1 dataset, with the iTransformer as the baseline method.
@@ V-C1 Effectiveness of $\mathcal{L}_{D}$ and $\rho$
[…]
    baseline with $\rho$, where the overall loss is $\mathcal{L}=\rho\mathcal{L}_{Y}$.
[…]
@@
    baseline with $\mathcal{L}_{D}$ and $\alpha$ (i.e., replacing $\rho$ by a learning weight $\alpha$), where the overall loss is $\mathcal{L}=\alpha\mathcal{L}_{Y}+(1-\alpha)\mathcal{L}_{D}$.
[…]
    baseline with TDAlign, where the overall loss is $\mathcal{L}=\rho\mathcal{L}_{Y}+(1-\rho)\mathcal{L}_{D}$.
[…]
V-C2 Comparison of the Formulation of Temporal Dependencies
As shown in Figure 4, we evaluate our choice of using first-order differencing values (i.e., change values between adjacent time steps) to formulate temporal dependencies by comparing it against three alternative formulations:
  • Using $\tau$-order differencing values to capture more complex relations. We experiment with $\tau\in\{1,2,3,4\}$.
  • Using change values between time steps separated by an interval $k$ to capture periodic patterns. We experiment with $k\in\{1,6,12,24,48\}$.
  • Using frequency domain features to capture global structures, as exemplified by the recent work FreDF [51]. As noted in the original work, FreDF necessitates manual tuning of the relative weight to balance the loss terms, with optimal values varying across datasets and prediction lengths. To ensure a fair comparison, we experiment with an exhaustive grid search over $\{0.1,0.2,\ldots,1.0\}$ and report its best-performing results, denoted as $\text{FreDF}_{best}$.
@@ V-C2 Comparison of the Formulation of Temporal Dependencies
The results in Figure 4(a) show that using first-order differencing values (i.e., $\tau=1$) achieves the best performance, with performance degrading as $\tau$ increases. Although higher-order differencing values represent more complex temporal dependencies, they may not provide substantial additional information compared to first-order differencing values. Instead, they tend to overshadow lower-order dependency information and over-constrain the model’s learning, thereby compromising its learning and generalization ability. Moreover, since each $d^{(\tau)}_{t+i}$ involves more time steps, higher-order differencing values become more sensitive to noise, leading to degraded model performance. The results in Figure 4(b) indicate that using change values between adjacent time steps (i.e., $k=1$) yields the best performance. This phenomenon can be attributed to the fact that as $k$ increases, the correlation between $x_{t+i}$ and $x_{t+i+k}$ typically weakens. Learning unreliable temporal dependency information adversely affects model performance. The results in Figure 4(b) further validate the effectiveness of our temporal dependency formulation. Although learning intricate correlations from the frequency domain yields substantial improvements, FreDF requires labor-intensive searching for the optimal relative weight. In contrast, our approach demonstrates superior performance while possessing greater flexibility in diverse forecasting scenarios. These results demonstrate that using first-order differencing values (i.e., change values between adjacent time steps) to formulate temporal dependencies is simple yet effective.
(a) ETTh1, $L=96$, $H=96$
(b) ETTh1, $L=96$, $H=192$
(c) ETTh2, $L=336$, $H=336$
(d) ETTh2, $L=336$, $H=720$
Figure 5: Robustness of TDAlign under Gaussian Noise with varying variances. (a)-(b) use iTransformer as the baseline method, (c)-(d) use DLinear as the baseline method.
(a) Evolution of $\mathcal{L}_{Y}$.
(b) Evolution of $\mathcal{L}_{D}$.
(c) Evolution of $\rho$.
(d) Validation $\text{MSE}_{D}$ and $\rho$.
Figure 6: Learning curves of TDAlign on the ETTh1 dataset, with iTransformer as the baseline method.
TABLE VIII: Computational cost of TDAlign on the ETTh1 dataset, with iTransformer as the baseline method.
$H$ Method Learnable params MACs Training time
$(K)$ $(MB)$ $(seconds/epoch)$
96 baseline 841.568 295.928 4.460
+TDAlign + 0 + 0 + 0.189
192 baseline 866.240 304.579 4.647
+TDAlign + 0 + 0 + 0.053
336 baseline 3379.024 1188.758 4.679
+TDAlign + 0 + 0 + 0.197
720 baseline 3576.016 1257.964 4.524
+TDAlign + 0 + 0 + 0.002
@@
Figure 7: Comparison of long-term time series forecasting results between our TDAlign and six baseline methods.
V-C3 Computational Cost
The computational cost of TDAlign is presented in Table VIII. As shown in the table, TDAlign maintains the same number of learnable parameters ($K$) and multiply-accumulate operations (MACs) as the baseline method across various time series lengths. The only overhead introduced by TDAlign is a slight increase in training time per epoch. These statistics demonstrate that TDAlign achieves the aforementioned substantial performance improvements while incurring only a negligible additional computational overhead, making it a highly efficient module that can be integrated with baseline methods for effective TDT modeling.
V-D Additional Experiments
V-D1 Effectiveness on LLM-based Methods
TABLE IX: Effectiveness of our TDAlign over different LLM-based baselines. Results are averaged over 3 runs, with better performance highlighted in bold.
Architecture LLM-based
Method TimesFM +TDAlign Time-LLM +TDAlign
Metric MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.536±0.021 0.496±0.004 0.511±0.013 0.466±0.005 0.407±0.001 0.426±0.003 0.401±0.006 0.414±0.005
192 0.559±0.021 0.508±0.002 0.540±0.011 0.484±0.004 0.447±0.001 0.454±0.001 0.423±0.011 0.429±0.006
336 0.563±0.020 0.512±0.006 0.544±0.011 0.491±0.004 0.462±0.010 0.467±0.007 0.436±0.005 0.444±0.003
720 0.539±0.028 0.516±0.014 0.529±0.014 0.502±0.005 0.511±0.024 0.506±0.015 0.452±0.010 0.478±0.008
ETTh2 96 0.358±0.031 0.397±0.021 0.323±0.008 0.360±0.003 0.304±0.007 0.363±0.004 0.294±0.014 0.355±0.010
192 0.421±0.030 0.433±0.020 0.380±0.016 0.401±0.001 0.363±0.005 0.393±0.003 0.342±0.003 0.384±0.004
336 0.429±0.017 0.448±0.014 0.404±0.007 0.423±0.003 0.381±0.006 0.412±0.003 0.362±0.008 0.399±0.004
720 0.476±0.035 0.475±0.023 0.443±0.010 0.462±0.009 0.414±0.005 0.447±0.005 0.389±0.005 0.430±0.006
Avg error 0.485 0.473 0.459 0.449 0.411 0.434 0.387 0.417
To evaluate the generalizability of TDAlign to Large Language Model (LLM) based methods, we apply it to two state-of-the-art forecasters: TimesFM (the 500M version) [55] and Time-LLM (the base version) [56]. We assess their performance under a full-shot fine-tuning setting for LTSF, repeating each experiment three times to ensure reliable results.
As presented in Table IX, TDAlign yields consistent and substantial performance gains for both models across multiple datasets and prediction horizons. This underscores the broad applicability of TDAlign as a plugin for diverse forecasting architectures, including large-scale pre-trained models.
V-D2 Robustness to Noise
To examine the robustness of TDAlign to noise, we inject Gaussian noise with varying variances into the training set as in [57]. We assess the performance using two baseline methods: iTransformer on the ETTh1 dataset with $H\in\{96,192\}$, and DLinear on the ETTh2 dataset with $H\in\{336,720\}$.
Figure 5 demonstrates that TDAlign consistently outperforms baseline methods across all experimental noise levels. Specifically, by learning local temporal dependencies and adaptively adjusting the relative weight between loss terms, TDAlign delivers a robust performance advantage, even under high-noise conditions. Simple models such as DLinear benefit more from the explicit modeling of TDT than more complex architectures like iTransformer. Additionally, we observe that TDAlign’s performance advantage over baseline methods becomes less pronounced as noise levels increase, reflecting the inherent challenges of learning reliable temporal patterns in noisy data. However, given that our tested noise levels already represent substantial perturbations rarely encountered in practical time series applications, TDAlign maintains its practical robustness and applicability in real-world scenarios.
V-E Visualization
V-E1 Training Process Analysis
To provide an intuitive understanding of the training process of TDAlign, Figure 6 illustrates its learning curves on the ETTh1 dataset, using iTransformer as the baseline method with input and prediction lengths of 96. Subfigures (a-c) depict the evolution of the conventional forecasting objective $\mathcal{L}_{Y}$, the TDT learning objective $\mathcal{L}_{D}$, and the adaptive weight $\rho$ within the first training epoch. Particularly, $\rho$ not only acts as an adaptive weight to balance the two objectives but also serves as a metric representing the sign inconsistency ratio. Subfigure (d) tracks the validation metric MSED and the value of $\rho$ across all epochs for three distinct experimental settings:
[…]
@@ V-E1 Training Process Analysis
[…]
    baseline with TDAlign, where the overall loss is $\mathcal{L}=\rho\mathcal{L}_{Y}+(1-\rho)\mathcal{L}_{D}$.
Figure 6 (a–c) reveals two key insights into the benefits of TDAlign. First, TDAlign significantly enhances TDT learning, as evidenced by its ability to achieve substantially lower $\mathcal{L}_{D}$ and $\rho$ values compared to the baseline method, while maintaining comparable primary forecasting performance ($\mathcal{L}_{Y}$). Second, the proposed dynamic weight $\rho$ effectively balances and steers the training process toward more efficient learning, as reflected in the faster convergence of both $\mathcal{L}_{D}$ and $\rho$ for TDAlign.
[…]
V-E2 Case Study
To illustrate the generalizability and effectiveness of TDAlign across datasets with diverse characteristics, Figure 7 visualizes forecasting results from each baseline method on two distinct datasets, covering all baseline methods and collectively spanning all datasets used in our main experiments. Overall, baseline methods produce inaccurate predictions with significant deviations from the ground truth. A closer examination reveals that these methods struggle to accurately capture both the change values and the change directions. Their predictions exhibit unrealistic fluctuations and jagged patterns, indicating their failure to learn the true temporal dynamics. This behavior is inconsistent with the inherent local smoothness of real-world time series, where consecutive observations transition gradually rather than abruptly. In contrast, TDAlign consistently generates more accurate predictions that closely align with the ground truth both locally and globally. Moreover, it mitigates the unrealistic fluctuations observed in baseline methods, resulting in smoother and more realistic predictions that better reflect the inherent local coherence of real-world time series. These improvements are primarily attributed to the proposed TDT learning objective, which explicitly captures variation patterns inherent in time series. Furthermore, by dynamically adjusting the importance of the TDT learning objective relative to the conventional forecasting objective based on the model’s learning state, TDAlign achieves a balanced optimization that enhances both local temporal modeling and global prediction accuracy. Consequently, TDAlign demonstrates significant improvements over baseline methods across diverse forecasting scenarios.
VI Conclusion